OpenStudy (anonymous):

How many atoms are in 2.56 mol of (NH4)3 PO4?

OpenStudy (anonymous):

\[2.56 \text{ mol} (NH_4)_3PO_4\times \frac{6.02 \times 10^{23} \text{ molecules}(NH_4)_3PO_4}{1 \text{ mol}(NH_4)_3PO_4} \times \frac{20 \text{ atoms}}{1 \text{ molecule}(NH_4)_3PO_4}\]

OpenStudy (anonymous):

First you should count the number of atoms in the molecule. There are 3N, (3x4)=12H, 1P and 4O. That's 3+12+1+4 = 20atoms in the molecule. 1 Mole of a substance is 6.02x10^23 molecules (avogadro's number) Therefore we have 20 x 6.02x10^23 x 2.56 = 3.08224x10^25 molecules